cnsider the following data rom wo populations are normally distributed. dependent samples th equal population var ances on struct a 9 % con der ce te val o es mate he dif rn ein po ulation means. Assume the popula on variances are equal and at the X1-36.42 S1-8.8 n1 #18 82 9.1 n2 19 ge 1 Click here to see the t distribution table page 2 The 90% confidence interval is( Round to two decimal places as needed.) DD
Answer #1

We will use t-statistic and two sample means (M1 and M2) to generate an interval estimate of the difference between two population means (μ1 and μ2).

The formula for estimation is:

μ1 - μ2 = (M1 - M2) ± ts(M1 - M2)

where:

M1 & M2 = sample means
t = t statistic determined by confidence level
s(M1 - M2) = standard error = √((s2p/n1) + (s2p/n2))

Detailed calculation is as shown below:


Pooled Variance
s2p = ((df1)(s21) + (df2)(s22)) / (df1 + df2) = 2807.06 / 35 = 80.2 where df1 = n1 - 1 = 17 and df2 = n2 - 1 = 18

Standard Error
s(M1 - M2) = √((s2p/n1) + (s2p/n2)) = √((80.2/18) + (80.2/19)) = 2.95

Confidence Interval
μ1 - μ2 = (M1 - M2) ± ts(M1 - M2) = 4.2 ± (1.69 * 2.95) = 4.2 ± 4.977

Hence,

μ1 - μ2 = (M1 - M2) = 4.2, 90% CI [-0.777, 9.177].

You can be 90% confident that the difference between your two population means (μ1 - μ2) lies between -0.777 and 9.177.